# Who You Are Matters: Bridging Topics and Social Roles via LLM-Enhanced Logical Recommendation

### v2

@@ B.1 Task Formulation
Top-$N$ Recommendation Task Define the set of users $\mathcal{U}$ and the set of items $\mathcal{I}$. The observed user-item interactions are represented as user histories $\mathcal{H}$, where each user’s history $\mathcal{H}_{u}\in\mathcal{I}^{n_{u}}$ has length $n_{u}$. For top-$N$ recommendation, the objective is to learn a scoring function $P(i|u)$ that suggests top-$N$ items (from $\mathcal{I}\setminus\mathcal{H}_{u}$) that each user $u$ is highly likely to engage with, and the ground truth positive target item set is denoted as $\mathcal{I}_{u}$. Following the collaborative filtering paradigm, we assume a binary interaction label $y_{u,i}$, indicating the user $u$’s positive feedback on an item $i$, and we also allow an additional reward weight signal $w_{i^{+}}$ for each positive item $i^{+}\in\mathcal{I}_{u}$ to accommodate multi-behavior scenarios. In terms of the model input, we focus on the user history modeling which is related to the tag-based encoder in our solution. Yet, we remind readers that there may exist other context features that include but are not limited to user profile features, time features, device, and network features. How these features may integrate the tag-logic information is out of the scope of this paper, but worth further investigation.
@@
User/Item Tag Identification Task As we have introduced in Section 2.1, $\mathcal{C}$ denote the set of all possible item (topic) tags (e.g., headphone) and $\mathcal{T}$ denote the set of all possible user (role) tags (e.g., symphonist). We assume that neither $\mathcal{C}$ nor $\mathcal{T}$ are known in advance, so we need an automatic inference framework to solve them. Recall that one of our focuses in this work is the user role identification task which formally finds a subset of user tags $\mathcal{T}_{u}\subset\mathcal{T}$ that describes a given user $u$. As discussed in Section 1, it is impractical to directly ask users to provide this information, but we can solve it by first figuring the user tags $\mathcal{T}_{i}\subset\mathcal{T}$ related to each item $i$, then learn a tag-based model to infer $\mathcal{T}_{u}$. This complements the conventional viewpoint that first associates the item topic tags $\mathcal{C}_{i}\subset\mathcal{C}$ to each item $i$, then predicts the item tag user profile $\mathcal{C}_{u}\subset\mathcal{C}$.
@@ B.3 Prompt Designs
B.4 Tag Extraction Algorithm
In section 2.1, we introduce the process of cover set reduction which aims to find a small subset of the full tag set that can cover a sufficient number of items while ensuring the semantic differences between tags. We present the algorithm in Alg.1 and the process runs on a daily basis. The process iteratively includes a new tag into the cover set, and each newly included tag maximizes the coverage on the uncovered items (line 6) until no less than $\tau=99\%$ of the items has at least one tag included in the cover set. Tags that have not been recalled by any item in the last $\mathbb{D}$ days will be removed (line 11), indicating an out-of-date tag. In practice, we observe that the cover set converges (less than 10 tag removed or added per day) after 30 days of updates. The statistics of the resulting tag sets and their cover sets are summarized in Table 6 in Appendix D.1.
@@ B.4 Tag Extraction Algorithm
B.5 Logic Reasoning Process
@@
B.6 Augmentation Model Specification
@@
$\mathbf{r}_{i}^{(t)}=\bm{\alpha}_{i}\mathbf{T}_{i},\quad\bm{\alpha}_{i}=\text{softmax}$ $\left(\mathbf{W}\mathbf{T}_{i}+\mathbf{b}\right)$ (11)
@@
Tag-Logic Exploration in Learning and Inference Augmentation: Without loss of generality, we explain the exploration strategy on user role tags in TagCF-ut as the extended description of Section 2.3.2 and 2.3.3, and the solution in TagCF-it is symmetric. We start by considering the initial tag set $\mathcal{T}(0)$ that focuses on improving utility (i.e., TagCF-util), where $\mathcal{T}(0)$ may represent $\mathcal{T}_{i}(0)$ for a given item (inferred from $P(t|i)$) or $\mathcal{T}_{u}(0)$ for a given user (inferred from $P(t|u)$). Then we can use the U2I logic graph to find logically related item topic tags as $\mathcal{C}(1)=\{c|\exists t\in\mathcal{T}(0),\text{s.t. }(t,c)\in\mathcal{E}^{\text{U2I}\ast}\}$. Note that we use the distilled model $\varphi$ to generate graphs, and the corresponding scores could be used as weights of the edges. In this case, it can also use a soft method that selects the tags with aggregated weights as $\mathcal{C}(1)=\{c|w_{c}>\delta\}$, where $w_{c}=\sum_{t\in\mathcal{T}(0),(t,c)\in\mathcal{E}^{\text{U2I}\ast}}P(c|t)$. Finally, we can obtain the final exploration tag set $\mathcal{T}(1)$ by applying I2U logic on $\mathcal{C}(1)$ as $\mathcal{T}(1)=\{t|\exists c\in\mathcal{C}(1),\text{s.t. }(c,t)\in\mathcal{E}^{\text{I2U}\ast}\}$. Again, the corresponding soft method gets $\mathcal{T}(1)=\{t|w_{t}>\delta\}$, where $w_{t}=\sum_{c\in\mathcal{C}(1),(c,t)\in\mathcal{E}^{\text{I2U}\ast}}w_{c}$. To better align the scale of weights in $\mathcal{T}(0)$ and $\mathcal{T}(1)$, we normalize the weights so that they sum up to one. For better illustration of this process, we further provide case studies of the difference between $T(0)$ and $T(1)$ in Appendix D.3.
@@
Appendix C Experimental Settings
C.1 Online Experiments
@@ C.2 Offline Experiments
Evaluation Protocol. We include common ranking accuracy indicators such as $N$ and MRR@$N$, as well as diversity metrics like ItemCoverage@$K$ and GiniIndex@$N$ (denoted as Cover@$N$ and Gini@$N$, respectively). In this paper, we observe $N\in\{10,20\}$. For each experiment across all models, we run training and evaluation for five rounds with different random seeds and report the average performance.
@@
Appendix D Observations and Analysis
D.1 Statistics of Tag Sets
Table 6 summarizes the statistics of the full tag set of $\mathcal{T}$ and $\mathcal{C}$, as well as the reduced cover set $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$. In practice, we find that tags follow an extremely skewed frequency distribution (Figure 11), indicating that not all tags are identically useful and expressive. While a few general tags may be retrieved by a large number of items, there also exists a large number of precise but unique tags that cannot cover a sufficient number of items. As illustrated in section 2.1, this motivates our design on the cover set reduction module. In practice, the open full set size tends to expand at a considerable rate even after the 30-day observation period, while the reduced cover set quickly converges in the first few days.
@@ D.1 Statistics of Tag Sets
type full size daily expansion cover set size
user tag 2,976,845 266,185 7,633
item tag 50,208,782 3,805,332 20,956
@@
D.2 Statistics of Logic Graph
@@
D.3 Case Study of Tag-Logic Exploration
@@
Appendix E Additional Evaluation Results
E.1 (M)LLM Evaluation with Human Experts
@@ Appendix E Additional Evaluation Results
Table 7: Generated tag omparison results against GPT-4o.
Test Set (G+S)/(B+S) (G+S)/(B+S) 95% CI Win-Tie Rate G/S/B Details
359 videos 0.92 [0.993,1.35] 59.88% 125/90/144

### v3

@@ A.1 Task Formulation
Top-NN Recommendation Task Define the set of users $\mathcal{U}$ and the set of items $\mathcal{I}$. The observed user-item interactions are represented as user histories $\mathcal{H}$, where each user’s history $\mathcal{H}_{u}\in\mathcal{I}^{n_{u}}$ has length $n_{u}$. For top-$N$ recommendation, the objective is to learn a scoring function $P(i|u)$ that suggests top-$N$ items (from $\mathcal{I}\setminus\mathcal{H}_{u}$) that each user $u$ is highly likely to engage with, and the ground truth positive target item set is denoted as $\mathcal{I}_{u}$. Following the collaborative filtering paradigm, we assume a binary interaction label $y_{u,i}$, indicating the user $u$’s positive feedback on an item $i$, and we also allow an additional reward weight signal $w_{i^{+}}$ for each positive item $i^{+}\in\mathcal{I}_{u}$ to accommodate multi-behavior scenarios. In terms of the model input, we focus on the user history modeling which is related to the tag-based encoder in our solution. Yet, we remind readers that there may exist other context features that include but are not limited to user profile features, time features, device, and network features. How these features may integrate the tag-logic information is out of the scope of this paper, but worth further investigation.
@@
User/Item Tag Identification Task As we have introduced in Section 3.1, $\mathcal{C}$ denotes the set of all possible item (topic) tags (e.g., headphone) and $\mathcal{T}$ denotes the set of all possible user (role) tags (e.g., symphonist). We assume that neither $\mathcal{C}$ nor $\mathcal{T}$ are known in advance, so we need an automatic inference framework to solve them. Recall that one of our focuses in this work is the user role identification task which formally finds a subset of user tags $\mathcal{T}_{u}\subset\mathcal{T}$ that describes a given user $u$. As discussed in Section 1, it is impractical to directly ask users to provide this information, but we can solve it by first figuring the user tags $\mathcal{T}_{i}\subset\mathcal{T}$ related to each item $i$, then learn a tag-based model to infer $\mathcal{T}_{u}$. This complements the conventional viewpoint that first associates the item topic tags $\mathcal{C}_{i}\subset\mathcal{C}$ to each item $i$, then predicts the item tag user profile $\mathcal{C}_{u}\subset\mathcal{C}$.
@@ A.3 Prompt Designs
A.4 Tag Extraction Algorithm
In section 3.1, we introduce the process of cover set reduction which aims to find a small subset of the full tag set that can cover a sufficient number of items while ensuring the semantic differences between tags. We present the algorithm in Alg.1 and the process runs on a daily basis. The process iteratively includes a new tag into the cover set, and each newly included tag maximizes the coverage on the uncovered items (line 6) until no less than $\tau=99\%$ of the items has at least one tag included in the cover set. Tags that have not been recalled by any item in the last $\mathbb{D}$ days will be removed (line 11), indicating an out-of-date tag. In practice, we observe that the cover set converges (less than 10 tag removed or added per day) after 30 days of updates. The statistics of the resulting tag sets and their cover sets are summarized in Table 6 in Appendix C.1.
@@ A.4 Tag Extraction Algorithm
A.5 Logic Reasoning Process
@@
A.6 Augmentation Model Specification
@@
$\mathbf{r}_{i}^{(t)}=\bm{\alpha}_{i}\mathbf{T}_{i},\qquad\bm{\alpha}_{i}=\text{softmax}$ $\left(\mathbf{W}\mathbf{T}_{i}+\mathbf{b}\right)$ (11)
@@
Tag-Logic Exploration in Learning and Inference Augmentation: Without loss of generality, we explain the exploration strategy on user role tags in TagCF-ut as the extended description of Section 3.3.2 and 3.3.3, and the solution in TagCF-it is symmetric. We start by considering the initial tag set $\mathcal{T}(0)$ that focuses on improving utility (i.e., TagCF-util), where $\mathcal{T}(0)$ may represent $\mathcal{T}_{i}(0)$ for a given item (inferred from $P(t|i)$) or $\mathcal{T}_{u}(0)$ for a given user (inferred from $P(t|u)$). Then we can use the U2I logic graph to find logically related item topic tags as $\mathcal{C}(1)=\{c|\exists t\in\mathcal{T}(0),\text{s.t. }(t,c)\in\mathcal{E}^{\text{U2I}\ast}\}$. Note that we use the distilled model $\varphi$ to generate graphs, and the corresponding scores could be used as weights of the edges. In this case, it can also use a soft method that selects the tags with aggregated weights as $\mathcal{C}(1)=\{c|w_{c}>\delta\}$, where $w_{c}=\sum_{t\in\mathcal{T}(0),(t,c)\in\mathcal{E}^{\text{U2I}\ast}}P(c|t)$. Finally, we can obtain the final exploration tag set $\mathcal{T}(1)$ by applying I2U logic on $\mathcal{C}(1)$ as $\mathcal{T}(1)=\{t|\exists c\in\mathcal{C}(1),\text{s.t. }(c,t)\in\mathcal{E}^{\text{I2U}\ast}\}$. Again, the corresponding soft method gets $\mathcal{T}(1)=\{t|w_{t}>\delta\}$, where $w_{t}=\sum_{c\in\mathcal{C}(1),(c,t)\in\mathcal{E}^{\text{I2U}\ast}}w_{c}$. To better align the scale of weights in $\mathcal{T}(0)$ and $\mathcal{T}(1)$, we normalize the weights so that they sum up to one. For better illustration of this process, we further provide case studies of the difference between $T(0)$ and $T(1)$ in Appendix C.3.
@@
Appendix B Experimental Settings
B.1 Online Experiments
@@ B.2 Offline Experiments
Evaluation Protocol. We include common ranking accuracy indicators such as NDCG@$N$ and MRR@$N$, as well as diversity metrics like ItemCoverage@$K$ and GiniIndex@$N$ (denoted as Cover@$N$ and Gini@$N$, respectively). In this paper, we observe $N\in\{10,20\}$. For each experiment across all models, we run training and evaluation for five rounds with different random seeds and report the average performance.
@@
Appendix C Observations and Analysis
C.1 Statistics of Tag Sets
Table 6 summarizes the statistics of the full tag set of $\mathcal{T}$ and $\mathcal{C}$, as well as the reduced cover set $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$. Based on the statistics provided in Table 6, we find that item tags generally have a shorter lifespan compared to user tags. While the full tag sets for both users and items continuously expand without removal, the daily expansions reveal key distinctions. Item tags exhibit a significantly higher daily expansion, indicating more frequent updates. In contrast, user tags have a much smaller daily expansion and have nearly converged in the cover set. This smaller set size with less frequent expansion suggests that user tags are more stable and have a longer lifespan, whereas the high update frequency of item tags points to a shorter lifespan. In practice, we also find that tags follow an extremely skewed frequency distribution (Figure 11), indicating that not all tags are identically useful and expressive. While a few general tags may be retrieved by a large number of items, there also exists a large number of precise but unique tags that cannot cover a sufficient number of items. As illustrated in section 3.1, this motivates our design of the cover set reduction module. In practice, the open full set size tends to expand at a considerable rate even after the 30-day observation period, while the reduced cover set quickly converges in the first few days.
@@ C.1 Statistics of Tag Sets
type full size daily expansion cover set size cover set daily expansion
user tag 2,976,845 200-300K 7,633 converged
item tag 50,208,782 3.5-4.0 million 20,956 hundreds
@@
C.2 Statistics of Logic Graph
@@
C.3 Case Study of Tag-Logic Exploration
@@
Appendix D Additional Evaluation Results
D.1 (M)LLM Evaluation with Human Experts
Table 7: Generated tag comparison results against GPT-4o.
Test Set (G+S)/(B+S) (G+S)/(B+S) 95% CI Win-Tie Rate G/S/B Details
359 videos 0.92 [0.993,1.35] 59.88% 125/90/144
@@ Appendix D Additional Evaluation Results
Note: Compared with tag extraction, we keep a higher tolerance for the factual accuracy of the generated logic graphs, as their primary objective is to facilitate user interest exploration. This goal prioritizes diversity and the stimulation of potential user interests over strict factual precision. Nevertheless, to objectively assess the quality of these graphs, we conducted a corresponding human evaluation study. The results on a test set of 3,220 videos are summarized in the table 9.
Table 9: Tag logic graph comparison results against GPT-4o.
Test Set (G+S)/(B+S) (G+S)/(B+S) 95% CI Win-Tie Rate G/S/B Details
3,220 videos 0.875 [0.955, 1.19] 52.3% 1237/500/1483
D.2 Different LLM Size & Complexity
Different LLM Size: To determine the optimal LLM size for our TagCF framework in an industrial setting, we conducted extensive experiments with LLMs of various parameter scales, including 0.5B, 1B, 7B, and 9B versions. The key findings from our scaling study are summarized in the table 10. Our parameter scaling experiments found that while smaller models (0.5B/1B) handle 93% of cases, the 7B model is crucial for the hardest 7%. The 9B model offered only a marginal +3% accuracy gain but with significantly higher latency. Thus, we employ a cost-effective cascade of smaller models for easy cases and the 7B model for hard samples, achieving an optimal balance.
Complexity: To optimize computational efficiency, our system performs tag extraction in a threshold-based manner, processing only new videos that exceed a predefined interaction count (e.g., 500 interactions). This selective approach ensures that resources are allocated to higher-impact content while maintaining tagging quality. Furthermore, model distillation is employed to enhance inference efficiency, enabling the distilled model to extend coverage to all videos cost-effectively. For online integration, we leverage a highly efficient key-value (KV) database, which allows for the retrieval of a video’s associated tag set in constant $\mathcal{O}(1)$ time complexity. Our online workflow is designed to facilitate parallel reading of tags and subsequent modeling computations. Crucially, once extracted, the tags are stored as immutable metadata permanently linked to the video. This persistent tag set serves all downstream recommendation tasks throughout the video’s lifecycle, significantly enhancing the overall performance and reusability within the system.
Table 10: Different LLM Size Results (7B as Baseline)
Model Accuracy Coverage Hard Case Relative Cost
0.5B / 1B -7% -7% $\times$ -69%
7B - - $\checkmark$ -
9B +3% +0% $\checkmark$ +37%
Appendix E Broader Impacts
Our work on enhancing recommender systems through LLM-enhanced user role identification and logical Recommendation has significant societal implications, both positive and negative. By incorporating user roles and behavioral logic, our framework enables more nuanced recommendations, better aligning with individual preferences and social contexts. This can enhance user engagement and satisfaction in applications such as e-commerce, content platforms, and educational tools. On the other hand, the framework may potentially provide new methodologies to social science by providing automatic and systematic solutions to discover user behavioral logic in the big data era.
However, despite the advancements offered by our method, it is essential to acknowledge potential drawbacks. If the system misinterprets user roles or behavioral logic, it could lead to irrelevant or harmful recommendations. Additionally, concerns regarding privacy and fairness arise due to the collection and analysis of user data for recommendations, necessitating careful consideration of ethical implications in its deployment. To this end, further complementary research on the solutions to mitigate these issues is necessary to achieve a benign and protective recommender system for users.